OpenStudy (anonymous):

which method should i use to solve the binomial 0=x^2+6x

OpenStudy (eyust707):

bring 6x to the other side of the equation

OpenStudy (anonymous):

i need to solve for graphing far to early in the morning and my brain is fried!!! i need to solve for x intercepts

OpenStudy (eyust707):

-6x=x^2

OpenStudy (eyust707):

-6=x

OpenStudy (eyust707):

do you see how i did that?

OpenStudy (anonymous):

x(x+6)=0 factor out x x=0 or x+6=0 two solutions x=0 or x=-6

OpenStudy (eyust707):

ditto that lol its been so long

OpenStudy (anonymous):

Check: x=0 0^2+6*0=0 true x=-6 (-6)^2+6*(-6)=0 36-36=0 true